OpenStudy (anonymous):

Find the non-extraneous solutions of the square root of the quantity of x plus 3 minus 4 equals x minus 1.

zepdrix (zepdrix):

That's hard to read in text. Is this the problem?\[\Large\rm \sqrt{x+3}-4=x-1\]

jagr2713 (jagr2713):

isnt there suppose to be an answer choice